OpenStudy (anonymous):

write an equation in slope intercept form of the line: (5, 1); m=1/3

OpenStudy (anonymous):

y=(1/3)x -(2/3)

OpenStudy (anonymous):

Can you explain how you did this? I am not good at math at all. :P

OpenStudy (anonymous):

Ok!

OpenStudy (hunus):

\[y=mx+b\] \[(1)=(\frac{ 1 }{ 3 })(5)+b\] \[1-\frac{5}{3}=b\] \[b = \frac{3}{3}-\frac{5}{3}=\frac{-2}{3}\] So y=mx+b \[y=\frac{x}{3}-\frac{2}{3}\]

OpenStudy (anonymous):

Okay, Thank you so much! I gave you a medal. c:

OpenStudy (anonymous):

Oh Interesting ,Hunus has used a similar method to solve it, I think that is great!!!!!!!!!!!!!!!